# A Search for AGN sources of the IceCube Diffuse Neutrino Flux

### v2

@@ 1 Introduction
We then go on to incorporate neutrino energy information, new in IceCube’s 10 year catalog, and perform an energy-dependent analysis with a likelihood formulation that includes the provided “energy proxy” of the neutrino, focusing on the 4LAC sources in the Northern sky using the approach outlined in [27]. We restrict this analysis to the Northern sky because the IceCube public data set is dominated by the atmospheric muon background at declination angles less than ${}^{\circ}$. This boundary of ${}^{\circ}$ is a result of the provided smearing matrices, which describe the spread in reconstructed energy, direction, and angular uncertainty expected after developing studying Monte Carlo simulations of the IceCube detector. These matrices are provided with a central neutrino declination bin that covers the sky from ${}^{\circ}$ to ${}^{\circ}$ and thus includes backgrounds from the Southern sky. Without a dedicated cosmic ray simulation to characterize the energy dependence of the muon background, the relationship between the energy proxy and the true neutrino energy are not strongly correlated in the Southern sky. Thus, we choose to only look at the Northern sky where this background is negligible. While this analysis only focuses on the Northern sky, the addition of information about the neutrino energy distribution increases the sensitivity of the search overall; previous work found that with energy information included, the total number of events needed for a 5 $\sigma$ significance is reduced by a factor of two [27]. The same study also outlines a process that can be repeated in the Southern sky with a more robust energy reconstruction simulation that characterizes the muon background.
@@ 2 Method Overview
where $P^{sig}_{i}$, $\epsilon^{sig}_{i}$ and $T^{sig}_{i}$ are the respective spatial, energy, and time components for the signal PDF, $\sigma_{i}$ is the uncertainty associated with each event, given in terms of angular resolution provided by the data set, $\gamma$ is the spectral index, $t_{i}$ is the time of the event, and $\phi_{s}$ is the flux of the source as a function of time. The background PDF is defined similarly by:
$\Large B_{i}=P^{bkg}_{i}(\sigma_{i},\vec{x_{i}}|\vec{x_{s}})\cdot\epsilon^{bkg}_{i}(E_{i},\delta_{i}|\gamma)\cdot T^{bkg}_{i}(t_{i},\phi_{s}),$ (2.3)
@@ 2.1 Catalog Descriptions
The Fermi 4LAC catalog is comprised of active galactic nuclei (AGN) sources detected in the 50 MeV to 1 TeV energy range over an eight year period from August 2008 to August 2016. Of the 2863 total AGNs, 2796 are blazars and 63 are non-blazars. The blazars can be further broken down into 658 flat spectrum radio quasars (FSRQs), 1067 BL Lacs, and 1071 “blazars of unknown origin.” We do not consider sources within ${}^{\circ}$ of the poles.
The MOJAVE XV catalog consists of observations in the 15 GHz band of 437 AGN over the course of 20 years, between 1996 and 2016. During this time, 5321 measurements were taken of the AGN sources, with most sources visited between 5 and 15 times. These time-dependent measurements give unique insight into whether a source is flaring, either relative to its baseline or compared to similar sources. In this study, we only consider MOJAVE data taken after 2008, to align with the operational window of the IceCube instrument.
@@ 3.1 Results
In Fig. 1, we show an all-sky map of the likelihood of a neutrino point source, $\sqrt{2\Delta\ln\mathcal{L}}$, in terms of right ascension and declination. The scan was performed in steps of ${}^{\circ}$, and at each point we show the value of $n_{s}$ that maximizes the test statistic defined in Eq. 3.2. In Fig. 2 we show the distribution of this test statistic across the sky. When excluding the points containing NGC 1068, a 4.2 $\sigma$ source discovery by IceCube [7], the distribution is consistent with a normal distribution. In Table 1, we present the most significant locations in the sky other than NGC 1068, along with the associated likelihoods and pre- and post-trials $p$-values. The post-trials $p$-values are calculated by scrambling the right ascension of the data to create 1000 randomly distributed background-like distributions, and calculating how often these background-like distributions are more extreme than the real data. There is no evidence of a statistically significant source.
@@
$2\Delta\ln\mathcal{L}$ Pre-Trials $p$-value Post-Trials $p$-value RA Dec
18.15 2.0 $\times$ $10^{-5}$ 0.451 12.60 -67.40
17.72 2.6 $\times$ $10^{-5}$ 0.453 296.40 -21.00
17.55 2.8 $\times$ $10^{-5}$ 0.455 300.20 -32.80
17.47 2.9 $\times$ $10^{-5}$ 0.456 327.00 -66.60
17.31 3.2 $\times$ $10^{-5}$ 0.456 232.80 1.40
Table 1: The five most significant independent locations found in our spatial-only all-sky scan, excluding the coordinates of NGC 1068. After accounting for an appropriate trials factor, none of these potential sources are statistically significant.
@@ 4.1 Results
$2\Delta\ln\mathcal{L}$ Pre-Trials $p$-value Post-Trials $p$-value RA Dec
@@
Table 2: The five most significant independent locations found in our energy-dependent all-sky scan of only the Northern Hemisphere. After accounting for an appropriate trials factor, none of these potential sources are statistically significant.
With the test statistic based on the new likelihood formulation in hand, we find the updated distribution of the test statistic shown in the right panel Fig. 4. This distribution falls along a normal distribution, consistent with no significant source excesses. The pre- and post-trial $p$-values are calculated in the same was as in Section 3.1. The most significant locations on the sky in this search are presented in Table 2. NGC 1068 falls just below the lower boundary of the analysis in declination and is thus excluded.
@@ 5 A Search for Correlations with MOJAVE Radio AGN
We perform the analysis using 436 radio-loud AGN from the MOJAVE catalog. Due to detector location and limits, the catalog only consists of sources higher than ${}^{\circ}$ in declination. Each source in the catalog has a minimum of 5 flux density recordings at different times. Two sources are eliminated from the analysis, MOJAVE source 0415+379 and source 2200+420, as all of their flux density recordings were above the 2 $\sigma$ threshold and can thus be categorized as always flaring. These two sources could be analyzed independently of this temporal analysis. This methodology was chosen to fairly assess all of the sources in the MOJAVE catalog, which can individually have as few as five data points and as many as 30 over the two decades of data.
@@
The time-dependent component of the background PDF is independent of seasonal variations [31, 36], but is dependent on the declination and year of the neutrino event. An additional time-dependence arises due to the construction of IceCube itself: more atmospheric muons were recorded in the first three years of this dataset than later years due to the phased construction of the detector. This prevalence of events with declination ${}^{\circ}$ during these early years is described in the background PDF, with unique PDFs defined as a function of declination for each 0.1 year bin. After the first 3 years, the PDF no longer changes with time, only declination.

### v3

@@ 1 Introduction
We then go on to incorporate neutrino energy information, new in IceCube’s 10 year catalog, and perform an energy-dependent analysis with a likelihood formulation that includes the provided “energy proxy” of the neutrino, focusing on the 4LAC sources in the Northern sky using the approach outlined in [27]. We restrict this analysis to the Northern sky, defined conservatively as events with a declination of > 10, due to the coarse binning of the smearing matrices provided by the IceCube public data set. These smearing matrices describe the spread in reconstructed energy, direction, and angular uncertainty expected after developing Monte Carlo simulations of the IceCube detector. In the Southern sky, the lack of a dedicated cosmic ray simulation makes characterizing the muon background too challenging, and thus this region is omitted from this analysis. We also choose to omit the horizon region, defined by the smearing matrices as -10 to 10, to be sure any edge effects in the detector response are not impacting our analysis. Thus, only the Northern sky above 10 declination remains in the analysis. Even still, the addition of information about the neutrino energy distribution increases the sensitivity of the search overall; previous work found that with energy information included, the total number of events needed for a 5 $\sigma$ significance is reduced by a factor of two [27]. The same study also outlines a process that can be repeated in the Southern sky with a more robust energy reconstruction simulation that characterizes the muon background.
@@ 2 Method Overview
where $P^{sig}_{i}$, $\epsilon^{sig}_{i}$ and $T^{sig}_{i}$ are the respective spatial, energy, and time components for the signal PDF, $\sigma_{i}$ is the uncertainty associated with each event, given in terms of angular resolution provided by the data set, $E_{i}$ is the energy proxy, $\gamma$ is the spectral index (set to either 2 or 2.5 for each analysis), $t_{i}$ is the time of the event, and $\phi_{s}$ is the flux of the source as a function of time. The background PDF is defined similarly by:
$\Large B_{i}=P^{bkg}_{i}(\delta_{i})\cdot\epsilon^{bkg}_{i}(E_{i})\cdot T^{bkg}_{i}(t_{i}),$ (2.3)
@@ 2.1 Catalog Descriptions
The Fermi 4LAC catalog is comprised of active galactic nuclei (AGN) sources detected in the 50 MeV to 1 TeV energy range over an eight year period from August 2008 to August 2016. Of the 2863 total AGNs, 2796 are blazars, 63 are non-blazars, and 4 are unidentified. The blazars can be further broken down into 658 flat spectrum radio quasars (FSRQs), 1067 BL Lacs, and 1071 “blazars of unknown origin.” We do not consider sources within 3 of the poles.
The MOJAVE XV catalog consists of observations in the 15 GHz band of 437 AGN over the course of 20 years, between 1996 and 2016. This list of sources includes 265 quasars, 127 BL Lacs, 27 Radio Galaxies, 5 Seyfert-1 Galaxies, and 13 unidentified objects. During this time, 5321 measurements were taken of the AGN sources, with most sources visited between 5 and 15 times. These time-dependent measurements give unique insight into whether a source is flaring, either relative to its baseline or compared to similar sources. In this study, we only consider MOJAVE data taken after 2008, to align with the operational window of the IceCube instrument.
@@ 3.1 Results
To validate our methodology, in Fig. 1 we show an all-sky map of the likelihood of a neutrino point source, $\sqrt{2\Delta\ln\mathcal{L}}$, in terms of right ascension and declination. The scan was performed in steps of 0.2, and at each point we show the value of $n_{s}$ that maximizes the test statistic defined in Eq. 3.2. In Fig. 2 we show the distribution of this test statistic across the sky. When excluding the points containing NGC 1068, a 4.2 $\sigma$ source discovery by IceCube [7], the distribution is consistent with a normal distribution. In Table 1, we present the most significant locations in the sky other than NGC 1068, along with the associated likelihoods and pre-trial $p$-values. The post-trial $p$-value for the most significant source is calculated by scrambling the right ascension of the data to create 1000 randomly distributed background-like distributions, and calculating how often these background-like distributions are more extreme than the real data. There is no evidence of a statistically significant source, which is the expected result, consistent with other all-sky scans of the IceCube dataset [33].
@@
$2\Delta\ln\mathcal{L}$ Pre-Trials $p$-value RA Dec
18.15 2.0 $\times$ $10^{-5}$ 12.60 -67.40
17.72 2.6 $\times$ $10^{-5}$ 296.40 -21.00
17.55 2.8 $\times$ $10^{-5}$ 300.20 -32.80
17.47 2.9 $\times$ $10^{-5}$ 327.00 -66.60
17.31 3.2 $\times$ $10^{-5}$ 232.80 1.40
Table 1: The five most significant independent locations found in our spatial-only all-sky scan, excluding the coordinates of NGC 1068. The most significant source has a post-trial $p$-value of 0.451, which is not statistically significant.
@@ 4.1 Results
$2\Delta\ln\mathcal{L}$ Pre-Trials $p$-value RA Dec
@@
Table 2: The five most significant independent locations found in our energy-dependent all-sky scan of only the Northern Hemisphere. The most significant source has a post-trial $p$-value of 0.456, which is not statistically significant.
With the test statistic based on the new likelihood formulation in hand, we find the updated distribution of the test statistic shown in the right panel Fig. 4. This distribution falls along a normal distribution, consistent with no significant source excesses. The pre- and post-trial $p$-values are calculated in the same was as in Section 3.1. The most significant locations on the sky in this search are presented in Table 2. While not statistically significant in this search, we do note that the brightest spot in our likelihood map, with a Right Ascension=182.60 and Declination=39.40, is within two arc minutes of NGC 4151, a source with recent evidence of neutrino emission [37]. NGC 1068 falls just below the lower boundary of the analysis in declination and is thus excluded.
@@ 5 A Search for Correlations with MOJAVE Radio AGN
We perform the analysis using 436 radio-loud AGN from the MOJAVE catalog. Due to detector location and limits, the catalog only consists of sources higher than -30 in declination. Each source in the catalog has a minimum of 5 flux density recordings at different times. Two sources are eliminated from the analysis, MOJAVE source 0415+379 and source 2200+420, as all of their flux density recordings were above the 2 $\sigma$ threshold and can thus be categorized as always flaring. These two sources could be analyzed independently of this temporal analysis. This methodology was chosen to fairly assess all of the sources in the MOJAVE catalog, which can individually have as few as five data points and as many as 30 over the two decades of data.
@@
The time-dependent component of the background PDF is independent of seasonal variations [31, 38], but is dependent on the declination and year of the neutrino event. An additional time-dependence arises due to the construction of IceCube itself: more atmospheric muons were recorded in the first three years of this dataset than later years due to the phased construction of the detector. This prevalence of events with declination <0 during these early years is described in the background PDF, with unique PDFs defined as a function of declination for each 0.1 year bin. After the first 3 years, the PDF no longer changes with time, only declination.